Friday, September 16, 2011

Night Log 2011-09-15

SA: Tim
SO: Siphelo


beautiful photometric night with steady 0.7"-1.2" seeing. fairly smooth operations for the most part. some RSS issues early on and a couple of payload failures, but they were all fairly quickly recovered from.

- no nightlog for 20110914 since everything was shut down in preparation for the UPS replacement. spent first half of night working on improving pointing model for the mass-dimm. high winds in the 2nd half forced early closure of that.

- UPS replacement successful.

- replaced BVRI with g', r', i', and z' in SALTICAM.

- removed P000160N13 and P000160N15 from RSS and replaced with P000277N06 and P000277N08 for Barger et al.

- got data for the following programs: 2011-3-UKSC_POL_RSA-001, 2010-1-RSA_UKSC_GU-001, and 2011-3-POL-005. got engineering data in the form of biases for both SALTICAM and PDET as well as some data to measure ADC performance as a function of elevation and prism separation.

No comments:

Post a Comment